# A Survey of Embodied AI in Healthcare: Techniques, Applications, and Opportunities

### v1

@@ 6.3.4 Stronger Generalization Capabilities
The future of EmAI in healthcare lies in its potential to seamlessly integrate physical and AI capabilities, transforming the way healthcare systems interact with patients and their environments. Currently, most research in this field has focused on discrete applications—such as robotic-assisted surgery or rehabilitation—yet the broader frontier is in developing EmAI for more generalized settings. To this end, the exploration of spatial intelligence, multi-modality foundational models, large language models, world models, and physical AI is crucial [352].
@@ 8 Acknowledgement
This research was supported by an internal fund from the Hong Kong University of Science and Technology (Guangzhou).

### v2

@@ 6.3.4 Stronger Generalization Capabilities
The future of EmAI in healthcare lies in its potential to seamlessly integrate physical and AI capabilities, transforming the way healthcare systems interact with patients and their environments. Currently, most research in this field has focused on discrete applications—such as robotic-assisted surgery or rehabilitation—yet the broader frontier is in developing EmAI for more generalized settings. To this end, the exploration of spatial intelligence, multi-modality foundational models, large language models, world models, and physical AI is crucial [352]. Additionally, evolutionary algorithms that autonomously tune hyperparameters and support continuous learning also help models to evolve and better adapt post-deployment in diverse environments.
@@ 8 Acknowledgement
This research was supported by the internal fund from the Hong Kong University of Science and Technology (Guangzhou).